# ARS: Automatic Routing Solver with Large Language Models

### v1

@@ Appendix G Database
The database contains six representative constraints, which exemplify problems related to these constraints. If no example problem is associated with the given input $I$, the base case $(I_{0},C_{0})$, which includes no additional constraints, will be selected. The following sections introduce these six representative constraints and the base case.
G.1 No Additional Constraints
Constraint Description: No additional constraints.
@@ 5 Args:
14
15 Returns:
@@
8 - "edge_weight": A 2D NumPy array of distances between nodes.
@@
8 - "edge_weight": A 2D NumPy array of distances between nodes.
@@
8 - "edge_weight": A 2D NumPy array of distances between nodes.
@@
8 - "edge_weight": A 2D NumPy array of distances between nodes.
@@
8 - "edge_weight": A 2D NumPy array of distances between nodes.
@@
8 - "edge_weight": A 2D NumPy array of distances between nodes.

### v3

@@ Appendix G Database
The database contains six representative constraints, which exemplify problems related to these constraints. If no example problem is associated with the given input $I$, the base case $(I_{0},C_{0})$, which includes no relevant constraints, will be selected. The following sections introduce these six representative constraints and the base case.
G.1 No Relevant Constraints
Constraint Description: No relevant constraints.
@@ 5 Args:
10 - "edge_weight": A 2D NumPy array of node distances.
@@
13 - routes (list): A list of routes, where each route is a list of node IDs (integers, excluding the depot node 0)."
14 Return:
@@
8 - "edge_weight": A 2D NumPy array of node distances.
@@
8 - "edge_weight": A 2D NumPy array of node distances.
@@
8 - "edge_weight": A 2D NumPy array of node distances.
@@
8 - "edge_weight": A 2D NumPy array of node distances.
@@
8 - "edge_weight": A 2D NumPy array of node distances.
@@
8 - "edge_weight": A 2D NumPy array of node distances.
@@ 14 Note:
Appendix H Potential Societal Impact
Automatic Routing Solver (ARS) has the potential to drive significant societal advancements by addressing complex real-world Vehicle Routing Problems (VRPs) in industries such as logistics, transportation, and healthcare. ARS leverages Large Language Model (LLM) agents to automate the design of constraint-aware heuristic solvers, offering flexibility and efficiency in solving diverse VRP scenarios.
ARS brings two main strengths to VRP solutions: 1) it dynamically adapts to diverse practical constraints, providing robust solutions without requiring extensive manual design, and 2) it introduces interpretability, enabling decision-makers to better understand and customize routing solutions for specific needs. These features make ARS a valuable tool for optimizing operations, reducing costs, and improving resource utilization across sectors. Additionally, the RoutBench benchmark ensures rigorous evaluation of ARS, further validating its real-world applicability.
However, ARS is not without challenges. Over-reliance on automated solvers may limit human oversight, and misinterpretation of results could lead to suboptimal decisions. Furthermore, the use of LLMs in ARS raises concerns about data security, as sensitive operational or constraint information could inadvertently be exposed. Addressing these risks will be crucial to ensuring ARS’s safe and ethical deployment. Lastly, while ARS demonstrates strong performance, its success rate may vary depending on the complexity of constraints, potentially delaying decision-making in highly intricate scenarios.
Appendix I Licenses
The licenses and URLs of the baseline methods are provided in Table 12.
Table 12: A summary of licenses.
Resources Type URL License
CPLEX Code https://www.ibm.com/products/ilog-cplex-optimization-studio Available for academic research use
OR-Tools Code https://github.com/google/or-tools Apache License 2.0
Gurobi Code https://www.gurobi.com/ Available for academic research use
LKH-3 Code http://webhotel4.ruc.dk/~keld/research/LKH-3/ Available for academic research use
PyVRP-HGS Code https://github.com/PyVRP/PyVRP MIT License
Reflexion Code https://github.com/noahshinn/reflexion MIT License
PHP Code https://github.com/chuanyang-Zheng/Progressive-Hint Available online
CoE Code https://github.com/xzymustbexzy/Chain-of-Experts/tree/main Available online
Self-verification Code https://github.com/Zhehui-Huang/LLM_Routing MIT License
Solomon Dataset http://vrp.atd-lab.inf.puc-rio.br/index.php/en/ Available for academic research use
CVRPLib Dataset http://vrp.atd-lab.inf.puc-rio.br/ Available for academic research use